40m QIL Cryo_Lab CTN SUS_Lab CAML OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 11513     Entry time: Tue Aug 18 03:56:09 2015
Author: ericq 
Type: Update 
Category: LSC 
Subject: locking efforts 

Now that the updated ALS is stable, and the PRC angular FF is revived, I've been working on relocking PRFPMI. While the RMS arm fluctuations are surely smaller than they used to be, there is no noticible difference to the ears when buzzing around resonance, but this doesn't really mean much. 

Frustratingly, I am not able to stably blend in any RF CARM error signal into the slow length control path (i.e. CARM_B). Bringing AS55 Q into DARM with the 20:0 integrator is working fine, but we really need to supress CARM to get anywhere. I'm not sure why this isn't working; poking around into the settings that were used when we were regularly locking didn't turn up any differences as far as I could tell. Investigations continue... 

Some minor changes to the locking script were made, to account for the increased ALS displacement sensitivity from the longer delay line. 


Since the ALS is now in a fairly stable state, I've updated the calibrated PSD template at /users/Templates/ALS/ALS_outOfLoop_Ref.xml, and added some coherence plots for some commonly coupled quantities (beat signal amplitude, IR error signal, green PDH error signal and green transmission). 

Attachment 1: newALSref.pdf  79 kB  Uploaded Tue Aug 18 05:03:47 2015  | Hide | Hide all | Show all
newALSref.pdf
Attachment 2: xCoh.pdf  139 kB  Uploaded Tue Aug 18 05:04:03 2015  | Show | Hide all | Show all
Attachment 3: yCoh.pdf  134 kB  Uploaded Tue Aug 18 05:04:51 2015  | Hide | Hide all | Show all
yCoh.pdf
ELOG V3.1.3-